OpenStudy (atlchic):

Solve by substitution, 3x+y=6 and y+2=x

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

you know that x = (y+2) from the second equation so you subsitute that into the first equation to get \[3(y+2) + y = 6\]

OpenStudy (anonymous):

this simplifies into \[3y + 6 +y = 6\] which you can simplify further into \[4y+6 = 6\]

OpenStudy (anonymous):

solving that equation, you get y=0 and then you plug that y into the second equation x= y+2 = 0+2 so x=2

OpenStudy (anonymous):

your solution is the point (2,0)

OpenStudy (atlchic):

Solve by elimination, x+8y=3 and 4x-2y=7

OpenStudy (anonymous):

ok well first you multiply the second equation by 4 all the way across \[16x-8y=28\]

OpenStudy (anonymous):

then you add the two equations together and get \[(16x+x) + (8y-8y) = (3+28)\]

OpenStudy (anonymous):

17x = 31 ==> x = 31/7

OpenStudy (anonymous):

then plug in and solve for y

OpenStudy (atlchic):

and next

OpenStudy (atlchic):

hello

OpenStudy (anonymous):

plus in the x value to the firs tequation (31/17) + 8y = 3 = 51/17

OpenStudy (anonymous):

8y = 20/17 y = (20/17)*(1/8) = 5/34